Please help!
I need to find an estimated volume for a field study I have just done and I'm not sure if I am doing it right.
I have a 500x500mm set square that is placed on the ground to measure the distribution of wrack (seaweed) within the square, I get an approximate covering of 40%. The depth of the wrack is measured and averages 4cm in depth. What is the volume in cm^3 within the square?
After doing this 6 times, the average percentage covering was 12.5% and the average depth was 1.25cm, do you just use the same formula as the first part with the 40% & 4cm? If not, what is the volume in cm^3 for this one too?

Answer by Theo(13342)

i'm not exactly sure what you trying to do, but my impression is this.
your area is 500 * 500 mm^2 which is equal to 250,000 mm^2.

40% of this area is covered with seaweed.

i would calculate that as .4 * 250,000 mm^2 = 100,000 mm^2 of the area is covered in seaweed.

you are told that the depth of the seaweed is an average of 4 cm.

since there are 10 mm for every cm, that means the average depth is 40 mm.

the volume of the seaweed would therefore be 100,000 * 40 = 4,000,000 mm^3.

if you are looking for the average depth of the seaweed over the whole 250,000 mm^2 area, then the formula would be:

250,000 * x = 4,000,000.

solve for x to get x = 4,000,000 / x = 16 mm which is equal to 1.6 cm.

the average depth of the seaweed over the whole 250,000 square mm is equal to 16 mm which is equal to 1.6 cm.